OpenStudy (taylev105):

If y = 4x – 1, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? {(2, 5), (4, 12), (9, 25)} {(–3, –10), (0, –2), (6, –18)} {(7, 24), (–1, –9), (–8, 35)} {(1, 3), (–2, –9), (0, –1)}d

OpenStudy (igreen):

It's asking which set of ordered pairs are true for y = 4x - 1.

OpenStudy (taylev105):

yes ok

OpenStudy (igreen):

Let's try this set: {(1, 3), (–2, –9), (0, –1)} Now let's first plug in (1, 3). y = 4x - 1 3 = 4(1) - 1 Simplify it, what do you get?

OpenStudy (igreen):

{(1, 3), (-2, -9), (0, -1)}

OpenStudy (taylev105):

omg im lost

OpenStudy (solomonzelman):

if you wanted to find if a point (6,2) is on the line y=4x-1 (for example) you would do this: 2=4(6)-1 2=24-1 2=23 so (6,2) is NOT on the line y=4x-1. If you wanted to find if a point (2,5) is on the line y=4x-1 (for example) you would do this: 5=4(2)-1 5=6-1 5=5 so (2,5) is on the line y=4x-1.

OpenStudy (anonymous):

pick first point of like (2,5).... then put the value of x (2) in the equation and see if gives 5 or not ....

OpenStudy (solomonzelman):

see how I am plugging in my points?

OpenStudy (taylev105):

yesss

OpenStudy (taylev105):

ok so

OpenStudy (solomonzelman):

so what answer do you pick? (or you want we can do it in a different form, of an arithmetic sequence, if you like...)

OpenStudy (taylev105):

im so confused lol

OpenStudy (solomonzelman):

look, y=4x-1 ..right? whatever the x is, we multiply it times 4 and subtract 1. so. when x=1, y=(1 * 4) - 1 ---> y=3 when x=2, y=(2 * 4) - 1 ---> y=8-1 ---> y=7 when x=3, y=(3 * 4) - 3 ---> y=12-1 ---> y=11 and on.... ~~~~~~~~~~~~~~~~~~~ or you can look at it as: when x =1, then y= 3, and when x=2, then y=7 So for each x, we add 4 for y. so the terms would come along: x = 1 2 3 y = 3 7 11 (when x=0, the y is: y=4(0)-1, y=-1 ) x = 0 -1 -2 y = -1 -5 -9

OpenStudy (solomonzelman):

I think I am offering you enough assistance to solve this.